# Disentangling Latent Shifts of In-Context Learning Through Self-Training

### v1

@@ Appendix B Lipschitz Continuity in Neural Networks
This property ensures that the function $f$ behaves smoothly, meaning small changes in the input lead to small changes in the output, which is crucial for robustness in neural networks, particularly for predictive models (Khromov & Singh, 2024).
@@ B.1 Relationship Between the Lipschitz Constant and the Jacobian Matrix
$[\mathbf{J}_{f}(\mathbf{x})]_{i,j}=\frac{\partial f_{i}(\mathbf{x})}{\partial x_{j}}.$
The spectral norm of the Jacobian matrix, denoted $\|\mathbf{J}_{f}(\mathbf{x})\|_{2}$, provides an upper bound on the Lipschitz constant $L$ (Latorre et al., 2020):
@@
The spectral norm represents the greatest possible rate of change in the function’s output for any input variation. However, calculating the exact spectral norm can be computationally expensive, especially for deep neural networks.
@@ B.2 Frobenius Norm as a Surrogate for the Lipschitz Constant
where $r$ is the rank of the matrix $\mathbf{A}$. The Frobenius norm provides an upper bound on the spectral norm and thus serves as a useful proxy for estimating the Lipschitz constant. This approximation is particularly useful in large-scale models, such as LLMs, where direct computation of the spectral norm is infeasible.
@@ B.3 Empirical Evaluation of Lipschitz Continuity
In our experiments, we approximate the Lipschitz constant by computing the Frobenius norm of example representations with respect to input embeddings. As shown in Figure 2(a), sticl demonstrates a significantly lower approximated Lipschitz constant compared to PBFT and ICL. This lower value suggests that sticl is more robust to input perturbations, which is a critical property for correcting pseudo-labels.
In summary, approximating the Lipschitz constant using the Frobenius norm provides an efficient tool for managing the smoothness of neural networks, enabling robust model behavior even in challenging scenarios (Sonoda & Murata, 2017; Khromov & Singh, 2024).
Appendix C Additional Results
@@
GLUE MMLU Model Method rte sst qnli mnli cola mrpc qqp math misc Llama 2 (7B) 0-shot 57.8 75.4 59.3 55.7 40.7 59.4 58.7 29.0 59.0 $n$-shot 69.24.3 89.82.1 74.25.9 63.32.8 54.33.5 66.92.4 64.71.5 37.54.8 80.05.3 PBFT 69.02.7 89.70.4 73.35.0 64.44.7 51.22.9 67.92.0 64.61.6 40.03.2 79.52.1 ICV 68.04.6 87.82.6 71.26.7 60.94.0 53.12.4 68.81.7 65.01.9 39.52.7 62.50.6 Batch-ICL 75.20.8 91.21.9 74.00.8 66.53.3 55.92.1 70.30.8 69.11.8 34.52.3 77.04.1 sticl-F 77.20.7 90.20.7 76.84.2 66.52.4 60.11.2 71.60.2 68.80.8 43.01.6 82.52.5 sticl-S 81.92.5 92.10.3 77.30.9 70.41.8 62.83.4 72.32.6 68.20.5 46.51.5 82.51.7 sticl-R 81.11.9 93.62.0 74.73.6 69.62.9 57.92.9 73.12.0 66.82.3 41.52.6 82.03.7
@@
GLUE MMLU Model $n$ Method rte sst qnli mnli cola mrpc qqp math misc Llama 3 (8B) $4$ $n$-shot 71.35.4 84.54.4 70.12.9 62.42.7 54.63.5 69.24.1 62.02.3 37.03.9 76.52.5 sticl-S 80.31.5 90.90.9 76.31.4 70.11.8 61.42.0 72.91.5 70.31.2 43.01.3 77.51.8 $8$ $n$-shot 72.72.1 89.42.6 73.52.5 64.73.1 55.82.8 71.22.4 64.32.9 37.01.3 77.52.1 sticl-S 82.11.1 93.21.0 78.31.3 72.21.6 63.71.8 73.91.3 72.10.4 47.50.5 84.01.4 $32$ $n$-shot 75.33.2 93.21.9 77.72.9 69.11.9 58.31.5 76.42.2 74.21.9 43.01.5 84.52.1 sticl-S 87.90.6 97.90.4 83.10.9 74.01.1 64.61.2 79.40.6 74.81.5 56.50.2 89.00.4
Table 7: ID generalization scores of sticl-S for $n=16$ shots and $|\mathcal{D}_{\text{unlab}}|=200,500$ for Llama 3 (8B). Results are shown for GLUE datasets with $n$-shot and sticl-S methods. The standard deviations of $10$ runs are shown as subscripts.
@@
Model Method qnli $\rightarrow$ rte rte $\rightarrow$ qnli qqp $\rightarrow$ mrpc mrpc $\rightarrow$ qqp Phi 3 (mini 4k) $n$-shot $64.3_{2.5}$ $67.2_{1.5}$ $63.7_{2.3}$ $59.4_{2.2}$ PBFT $64.1_{1.8}$ $66.9_{1.6}$ $64.7_{2.0}$ $60.1_{1.4}$ sticl-S $67.4_{0.6}$ $69.2_{0.9}$ $66.3_{2.4}$ $64.4_{1.3}$ Llama 2 (7B) $n$-shot $62.9_{2.3}$ $66.3_{1.2}$ $64.5_{1.9}$ $61.1_{2.2}$ PBFT $62.8_{1.3}$ $68.1_{1.4}$ $65.9_{1.8}$ $61.3_{1.2}$ sticl-S $64.8_{0.4}$ $70.3_{0.6}$ $67.8_{2.1}$ $65.0_{1.1}$
Appendix D Experimental Details
D.1 Models
@@
D.2 Hyperparameters
We employ the AdamW optimizer (Loshchilov & Hutter, 2019) for both PBFT and sticl variants, with a learning rate of $10^{-4}$. For ICV (Liu et al., 2023) and Batch-ICL (Zhang et al., 2024), we follow the implementations provided in the original papers and adapt them to our codebase, using their default parameters where specified. In the case of Batch-ICL, we utilize attention heads from the last 20 layers ($k=20$) and fine-tune the model for $10$ epochs.
@@ D.2 Hyperparameters
D.3 Computing Infrastructure
We conducted our experiments on 4×4$\times$4 × AMD Ryzen Threadripper 3970X 32-Core Processors and 4×4$\times$4 × NVIDIA GeForce RTX 3090 GPUs with $24$GB of RAM.
@@ D.3 Computing Infrastructure
Appendix E Prompt Templates
E.1 GLUE Prompt Structure
@@
E.2 MMLU prompt structure
MMLU tasks are structured to provide the model with a set of demonstrations followed by a query. These prompts focus on multi-choice questions, where the model must predict one of the given options. Below is the general structure of the MMLU prompt template:

### v2

@@ Appendix B Lipschitz Continuity in Neural Networks
This property ensures that the function $f$ behaves smoothly, meaning small changes in the input lead to small changes in the output, which is crucial for robustness in neural networks, particularly for predictive models [20].
@@ B.1 Relationship Between the Lipschitz Constant and the Jacobian Matrix
$\mathbf{J}_{f}(\mathbf{x})]_{i,j}=\frac{\partial f_{i}(\mathbf{x})}{\partial x_{j}}.$
The spectral norm of the Jacobian matrix, denoted $\|\mathbf{J}_{f}(\mathbf{x})\|_{2}$, provides a pointwise lower bound on the global Lipschitz constant $L$:
@@
The spectral norm represents the greatest possible rate of change in the function’s output for any input variation. However, calculating the exact spectral norm can be computationally expensive, especially for deep neural networks, so the Frobenius norm is often used as an efficient alternative.
@@ B.2 Frobenius Norm as a Surrogate for the Lipschitz Constant
where $r$ is the rank of the matrix $\mathbf{A}$. While the Frobenius norm generally overestimates the spectral norm, the bounded gap between them implies that reducing the Frobenius norm below its initial value often corresponds to a decrease in the spectral norm as well. This relationship supports its use as a practical proxy for Lipschitz behavior: reductions in the Frobenius norm are generally associated with a lower Lipschitz constant, supported by evidence showing that the Lipschitz constant of neural networks tends to closely track the lower bound defined by the spectral norm [23, 20].
@@ B.3 Empirical Evaluation of Lipschitz Continuity
In our experiments, we approximate the Lipschitz constant by computing the Frobenius norm of the input–output Jacobian matrix, where the embeddings are the inputs and the penultimate layer produces the outputs. As shown in Figure 2(a), wilda demonstrates a significantly lower approximated Lipschitz constant compared to PBFT and ICL. This lower value suggests that wilda is more robust to input perturbations, which is a critical property for correcting pseudo-labels.
Appendix C Limitations
Computational cost.
wilda introduces additional computational overhead due to the fine-tuning of adapters. While this fine-tuning is more lightweight compared to full model fine-tuning, it remains more expensive than standard ICL, which avoids weight updates entirely. However, wilda offsets some of this cost by removing demonstrations from the input during inference. For instance, with Llama 3 (8B) processing $16$ demonstrations from GLUE datasets, inference takes approximately $120$ times longer than a 0-shot setup (processing only the query). This increased cost scales quadratically with the number of tokens, highlighting the self-attention mechanism as the primary bottleneck when handling $16$ demonstrations. Based on our measurements, fine-tuning with $100$ unlabeled instances and $16$ demonstrations using a single adapter corresponds to the computational cost of approximately $2100$ inferences in a $16$-shot setup. This implies that after about $2100$ inferences, the time spent on fine-tuning is effectively balanced by the reduction in per-inference computational cost.
Applicability.
wilda may be less suitable for scenarios with extremely limited resources, as it relies on access to a supply of unlabeled data. In our experiments with $\{4,8,16,32\}$ demonstrations, we typically used $100$ unlabeled instances, which proved sufficient to achieve strong performance. While unlabeled data is generally easier to acquire than labeled data, there may be scenarios where obtaining even a modest amount of unlabeled data is challenging, potentially limiting the applicability of wilda.
Large demonstration sets.
Although wilda efficiently encodes demonstrations into adapters to overcome context length limitations, the method has not been extensively tested with very large demonstration sets. From our findings, as the total number of demonstrations increases, using multiple adapters with manageable demonstration sizes tends to be more effective. For instance, we successfully employed $8$ adapters with $16$ demonstrations each (totaling $128$ demonstrations). While this approach theoretically allows for an indefinite increase in the number of demonstrations, its effectiveness with significantly larger sets remains unexplored. Moreover, using additional adapters increases computational costs, introducing a tradeoff between scalability and efficiency.
Appendix D Additional Results
D.1 Supplementary Tables
@@
GLUE MMLU Model Method rte sst qnli mnli cola mrpc qqp math misc Llama 2 (7B) 0-shot 57.8 75.4 59.3 55.7 40.7 59.4 58.7 29.0 59.0 $n$-shot 69.24.3 89.82.1 74.25.9 63.32.8 54.33.5 66.92.4 64.71.5 37.54.8 80.05.3 PBFT 69.02.7 89.70.4 73.35.0 64.44.7 51.22.9 67.92.0 64.61.6 40.03.2 79.52.1 ICV 68.04.6 87.82.6 71.26.7 60.94.0 53.12.4 68.81.7 65.01.9 39.52.7 62.50.6 Batch-ICL 75.20.8 91.21.9 74.00.8 66.53.3 55.92.1 70.30.8 69.11.8 34.52.3 77.04.1 wilda-F 77.20.7 90.20.7 76.84.2 66.52.4 60.11.2 71.60.2 68.80.8 43.01.6 82.52.5 wilda-S 81.92.5 92.10.3 77.30.9 70.41.8 62.83.4 72.32.6 68.20.5 46.51.5 82.51.7 wilda-R 81.11.9 93.62.0 74.73.6 69.62.9 57.92.9 73.12.0 66.82.3 41.52.6 82.03.7
@@
GLUE MMLU Model $n$ Method rte sst qnli mnli cola mrpc qqp math misc Llama 3 (8B) 44 $n$-shot 71.35.4 84.54.4 70.12.9 62.42.7 54.63.5 69.24.1 62.02.3 37.03.9 76.52.5 wilda-S 80.31.5 90.90.9 76.31.4 70.11.8 61.42.0 72.91.5 70.31.2 43.01.3 77.51.8 88 $n$-shot 72.72.1 89.42.6 73.52.5 64.73.1 55.82.8 71.22.4 64.32.9 37.01.3 77.52.1 wilda-S 82.11.1 93.21.0 78.31.3 72.21.6 63.71.8 73.91.3 72.10.4 47.50.5 84.01.4 $32$ $n$-shot 75.33.2 93.21.9 77.72.9 69.11.9 58.31.5 76.42.2 74.21.9 43.01.5 84.52.1 wilda-S 87.90.6 97.90.4 83.10.9 74.01.1 64.61.2 79.40.6 74.81.5 56.50.2 89.00.4
Table 7: ID generalization scores of wilda-S for $n=16$ shots and $|\mathcal{D}_{\text{unlab}}|=200,500$ for Llama 3 (8B). Results are shown for GLUE datasets with $n$-shot and wilda-S methods. The standard deviations of $10$ runs are shown as subscripts.
@@
Model Method qnli $\rightarrow$ rte rte $\rightarrow$ qnli qqp $\rightarrow$ mrpc mrpc $\rightarrow$ qqp Phi 3 (mini 4k) $n$-shot $64.3_{2.5}$ $67.2_{1.5}$ $63.7_{2.3}$ $59.4_{2.2}$ PBFT $4.1_{1.8} 66$$6.9_{1.6}$ $64.7_{2.0}$ $60.1_{1.4}$ wilda-S $67.4_{0.6}$ $9.2_{0.9} 66$$6.3_{2.4}$ $64.4_{1.3}$ Llama 2 (7B) $n$-shot $2.9_{2.3} 66$$6.3_{1.2}$ $64.5_{1.9}$ $61.1_{2.2}$ PBFT $62.8_{1.3}$ $68.1_{1.4}$ $65.9_{1.8}$ $61.3_{1.2}$ wilda-S $64.8_{0.4}$ $70.3_{0.6}$ $67.8_{2.1}$ $65.0_{1.1}$
D.2 Few-shot wilda
wilda is primarily designed for 0-shot operation, leveraging unlabeled data to encode task-specific information within the adapter. To examine its performance in few-shot setups, we evaluated wilda-S using Llama 3 (8B) in a $16$-shot configuration, where $16$ additional demonstrations were encoded into the adapter, resulting in a total of $32$ labeled instances. This setup was compared against standard $32$-shot ICL, as well as two wilda-S variants utilizing $32$ labeled instances in a 0-shot configuration. Additionally, we included a baseline for a 0-shot setup with only $16$ encoded demonstrations.
To standardize comparisons, we denote each wilda variant using the format $n/d$, where $n$ represents the number of shots (n-shot) and $d$ indicates the number of demonstrations encoded in the adapter. The results, averaged over 10 runs, are shown in Table 9.
Table 9: Performance comparison of wilda-S configurations and standard 32-shot ICL averaged over $10$ runs.
Method rte sst qnli mnli cola mrpc qqp math misc 32-shot ICL 75.3 93.2 77.7 69.1 58.3 76.4 74.2 43.0 84.5 wilda-S (0/32) 87.9 97.9 83.1 74.0 64.6 79.4 74.8 56.5 89.0 wilda-S (0/16) 86.0 96.1 81.4 73.1 64.3 77.7 73.1 49.5 88.0 wilda-S (16/16) 87.3 96.4 82.2 74.6 65.4 78.2 74.5 51.0 89.0
The results demonstrate that wilda-S in the $16$-shot configuration with $16$ encoded demonstrations ($16/16$) outperforms both standard $32$-shot ICL and wilda-S ($0/16$) across all datasets, showcasing its ability to utilize additional context during inference. However, it slightly underperforms compared to the 0-shot wilda-S variant with 32 encoded demonstrations ($0/32$), likely due to the fine-tuning process that is exclusive to the 0-shot setup. Nevertheless, the strong performance in n-shot setups ($n>0$) highlights the flexibility and efficacy of wilda-S in leveraging additional context provided within the prompt.
D.3 Faithful Encoding and Retrieval of Demonstrations
To evaluate whether demonstrations are faithfully encoded and disentangled, we conducted an experiment by encoding a single demonstration into the adapter and assessing the student model’s ability to capture this information. Specifically, we utilized $1000$ examples per dataset across the GLUE benchmark using Llama 3 (8B).
For each dataset, the student model was prompted with a simple instruction: “Repeat the demonstration word for word.” During the fine-tuning phase, the teacher model processed input examples using the following template: “Demonstration: {demonstration}. Answer: ({answer}).” The adapter learned to encode demonstration-specific information indirectly by aligning its outputs with the teacher’s responses, without explicitly seeing the demonstration itself. After training, the similarity between the student model’s response and the original demonstration was computed. Table 10 shows the average BERTScore similarity [42] between the original demonstrations and the student’s reconstructed response.
Table 10: Average BERTScore ($F_{1}$) similarity across GLUE datasets. Higher scores indicate better fidelity in recalling the encoded demonstration.
rte sst qnli mnli cola mrpc qqp
BERTScore 0.84 0.91 0.80 0.83 0.86 0.82 0.81
The consistently high BERTScore values across all datasets indicate that the student model can reliably retrieve the encoded demonstration from the adapter. This suggests that wilda effectively disentangles and stores task-specific information within the adapter’s weights. Notably, when compared to standard ICL, wilda often produced different outputs for certain queries, particularly in instances where it corrected “corrupted” labels provided by the teacher. Despite these differences, the student model maintained a high degree of semantic similarity in reproducing the demonstrations. This suggests that the adapter weights capture not only the demonstration itself but also additional latent information that contributes to improved generalization.
We present below a pair of examples from sst and rte, chosen to represent reconstructed demonstrations with similarity scores close to the dataset averages.
sst: Example 1 Original: Proves once again he hasn’t lost his touch, delivering a superb performance in an admittedly middling film. Answer: (Positive) Reconstructed: He demonstrates once more that he hasn’t missed a beat, delivering a remarkable performance in what is admittedly an average film. Answer: (Positive)
sst: Example 2 Original: Though many of the actors spark briefly when they first appear, they can’t generate enough heat in this cold vacuum of a comedy to ignite a reaction. Answer: (Negative) Reconstructed: Although some actors manage to show a hint of energy early on, they fail to create any real warmth or spark within this lifeless and chilly comedy. Answer: (Negative)
rte: Example 1 Original: Premise: The source added that the investigation proved that the bases of the genocide crime “were completed with a series of illegal arrests followed in some cases with assassinations or cases of disappearances and were preceded, according to information attached to the file, by cases of torture.” Hypothesis: Investigators discovered that a series of illicit arrests were often followed by disappearances or murders and were preceded by torture. Answer: (True) Reconstructed: Premise: The investigation confirmed that genocide involved illegal arrests followed by disappearances or murders, often preceded by torture. Hypothesis: Investigators found that unlawful arrests frequently resulted in disappearances or murders, often preceded by acts of torture. Answer: (True)
rte: Example 2 Original: Premise: American tobacco companies were showing a profit most quarters due to export sales of cigarettes and diversification of products sold, including food. Hypothesis: PM often entered markets with both cigarettes and food. Answer: (False) Reconstructed: Premise: Profitability was often maintained by American tobacco companies through diversification into food products and successful cigarette exports. Hypothesis: Philip Morris International offered food items and cigarettes. Answer: (False)
Appendix E Experimental Details
E.1 Models
@@
E.2 Hyperparameters
We employ the AdamW optimizer [28] for both PBFT and wilda variants, with a learning rate of $10^{-4}$. For ICV [27] and Batch-ICL [41], we follow the implementations provided in the original papers and adapt them to our codebase, using their default parameters where specified. In the case of Batch-ICL, we utilize attention heads from the last 20 layers ($k=20$) and fine-tune the model for $10$ epochs.
@@ E.2 Hyperparameters
E.3 Computing Infrastructure
We conducted our experiments on AMD Ryzen Threadripper 3970X 32-Core Processors and $4\times$ NVIDIA GeForce RTX 3090 GPUs with 2424GB of RAM.
@@ E.3 Computing Infrastructure
Appendix F Prompt Templates
F.1 GLUE Prompt Structure
@@
F.2 MMLU prompt structure